OpenStudy (anonymous):

a scientist begins with 100 mg of radioactive substance. after 6 days, it has decayed 60 mg. how long after the process began will it take to decay to 10mg

OpenStudy (campbell_st):

ok... so its exponential decay..?

OpenStudy (campbell_st):

\[A = A_{0}e^{-kt}\] A is the current size, Ao is the initial, k is the decay constant and t = time..

OpenStudy (campbell_st):

so the 1st step is to find the decay constant A = 60, Ao = 100 and t = 6 \[60 = 100 \times e^{-6k}\] you need to solve for k... before doing part 2.

OpenStudy (anonymous):

so how do u solve for k

OpenStudy (campbell_st):

ok so divide both sides by 100 \[0.6 = e^{-6k}\] take the base e log of both sides of the equation \[\ln(0.6) = -6k\] that should help you to find k. the 2nd part asks you to find t when A = 10... so a similar process

OpenStudy (anonymous):

would i take the ln of both sides or just 0.6

OpenStudy (campbell_st):

yes because \[\ln(e^a) = a\]

OpenStudy (anonymous):

woiuld k be=0.085

OpenStudy (campbell_st):

yep... you may need to go to around 4 or 5 decimal places as it will affect the answer in the next part. so same process to find t when A = 10

OpenStudy (anonymous):

what do u mean A? whats formula

OpenStudy (campbell_st):

ok... so you will have \[10 = 100 e^{-0.085138t}\]

OpenStudy (campbell_st):

now solve for t... similar process to the previous part.

OpenStudy (anonymous):

t=-27.047

OpenStudy (campbell_st):

time can't be negative..

OpenStudy (campbell_st):

I think you'll find its a positive value

OpenStudy (anonymous):

i keep getting a negative answer: 10=100e^-0.085138t i dont know how to solve it

OpenStudy (campbell_st):

ln(0.1) = -0.08513t t = ln(0.1)/-0.8513 will give a positive since the ln(0.1) is a negative